OpenStudy (anonymous):

Find the equation of the perpendicular AND parallel line to each of the following that passes through the point indicated -4/5x+y=-9 passing through the point (8,-3)

OpenStudy (cwrw238):

rewrite y = (4/5) x - 9 the parallel line will have the same slope - that is 4/5 use the general equation y-y1 = m(x-x1) where m = slope 4/5 and (x1,y1) = (8,,-3)

OpenStudy (cwrw238):

can yu do these substitutions?

OpenStudy (anonymous):

i need to find both parallel and perpendicular ,

OpenStudy (anonymous):

and the equation is -4/5x-9

OpenStudy (cwrw238):

yes ok - the perpendicular slope is - 1 / m where m is the slope. so in this case m = -1 / (4/5) = -5/4 so to find this equation substitute m = -5/4, x1 = 8 and y1 = -3 in the above equation

OpenStudy (cwrw238):

which is the correct equation - the one in the question or the one in your last post?

OpenStudy (anonymous):

Last post.

OpenStudy (anonymous):

i need the answers.

OpenStudy (cwrw238):

in that case the slope of the parallel line is -4/5 and of perpendicular line its -1 / (-4/5) = 5/4.

OpenStudy (cwrw238):

parallel line: y - (-3) = (-4/5)(x - 8) which can be simplified

OpenStudy (cwrw238):

perpendicular line: y - (-3) = (4/5)(x - 8)